Telescope Bootstrap Startup: 9 September 2012 UT

Observer: JHill (LBTO)
Telescope Operator: DHuerta (LBTO)
Software Support: CBiddick (Tucson)
AO Support: DMiller (Tucson)
Telescope Support: KNewton (LBTO)
Instrument Support: JMorris (LBTO)

Summary

Heavy Clouds at Sunset. This night has also been lost to weather - we left the retroreflector on the SX side. (It's a trick to get the sky to clear.) We briefly thought about opening - which is a significant improvement over the previous 2 nights.

We found a couple problems with 2012F software that need to be fixed.

Preparations

It took all day, but we tracked down the bad fiber, so AGW2 is properly on the network again.

Still running Build 2012B for the first half of the night.

Details

03:20 We changed a panel in DX AdSec Electronics Rack, and closed the door.

05:55 Start mirror ventilation. Humidity is down to 81%. Sky is still mostly cloudy.

06:00 Open 3x0.25 vent doors Chamber=17 T=9 D=6. Wind 5 m/sec from SW, but there is a thunderstorm down by Douglas. Left side vent door is not functional.

06:20 Sky is clearing. Humidity steady about 82%. Still a few drips coming off the dome.

Right MV duct pressure is 36, and left MV duct pressure is 38. MV seems to be operating normally.

06:34 Authorize for IRTC at RFBG (of course, IRTC is really mounted on the other side) 13107: Arbitrator interface generic error
The DX tertiary moved correctly.

06:43 TRACK preset - failed for preset flat command failing.

06:47 Authorize LUCI @ RFBG - preset flat failed.

06:51 Authorize LBTI on Both sides (M3s are starting at LFBG/RFBG)
DX M3 moved. SX M3 did NOT move, although the brake was disengaged for a while.

Switch TCS from 2012G to 2012F

07:00 Switch the TCS Build from 2012G to 2012F. That's what we should use for initial checkout even though 2012G has a good track record going. And we want to see how the tertiaries behave during 2012F authorizations. Humidity is back up to 88%, and vent doors have been closed.

07:14 Authorize IRTC @ RFBG - succeeds

07:19 TRACK preset - succeeds.

07:20 Humidty is up to 93% and we are starting fog/cloud over.

07:22 GUIDE preset (w/o fake star)

07:25 enable fake guide stars in gcs.conf, and restart GCS

07:26 GUIDE preset (w/ fake star) - Telescope off source too long, Why? Sep 9 07:27:56 tcs2 LBT_PCS: Side: right Centroid Xmm: 0.146845490093366 Ymm: 17.923760280663185 Xarcs: 0.245411707875360 Yarcs: 29.954618417149646

07:34 GUIDE preset again Telescope off source too long, This is Issue 4197.

Also notice that the guider images are not appearing in /Repository/20120909/GCS/ - this is Issue 4196.

07:53 ACQUIRE preset - no image appears

07:55 Authorize as LBTI on both sides (M3s starting at LFBG/RFBG)
Sun Sep 9 07:55:01.894 2012 TCS authorize command started for left LBTI at bentGregorianCenter focus and right LBTI at bentGregorianCenter focus
Sun Sep 9 07:56:20.622 2012 TCS authorize alert for left LBTI at bentGregorianCenter focus and right LBTI at bentGregorianCenter focus: Stop[2] complete
Sun Sep 9 07:56:20.673 2012 TCS authorize alert for left LBTI at bentGregorianCenter focus and right LBTI at bentGregorianCenter focus: Stop[12] complete
Sun Sep 9 07:56:20.724 2012 TCS authorize alert for left LBTI at bentGregorianCenter focus and right LBTI at bentGregorianCenter focus: sx.terc: [moveFocalStation.moveSelAbs.sendString] 
Sun Sep 9 07:56:20.774 2012 TCS authorize alert for left LBTI at bentGregorianCenter focus and right LBTI at bentGregorianCenter focus: PresetFlat[13] complete
Sun Sep 9 07:56:20.825 2012 TCS authorize command failed for left LBTI at bentGregorianCenter focus and right LBTI at bentGregorianCenter focus

This failed in the same manner as it did for 2012G an hour ago, except that the SX brake stayed disengaged for long after the authorize failed.

08:00 De-Init and Init the SX Tertiary as the brake never engaged.

08:03 Authorize LBTI on both sides (M3s already at the LBTI position) - succeeded.

08:05 Authorize LUCI on both sides (M3s starting at LBTI position) - IIF reported success, but SX M3 did not move.

08:09 Authorize LUCI on left only (SX M3 moves as expected) - success

08:11 Authorize LBTI on left only (SX M3 moves as expected) - success

08:14 Authorize LBTI on both sides (SX was at LBTI, DX was at RFBG) - IIF reported success, but DX M3 did not move.

08:26 Authorize LBTI on both sides (SX was at LBTI, DX was at RFBG) - this time we ended with both mirrors in the correct position.

10:10 Calling it a night with 98% humidty and fog.

-- JohnHill - 09 Sep 2012
Topic revision: r4 - 09 Sep 2012, JohnHill
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback